OpenStudy (anonymous):

I anyone up to helping me to solve a quadratic multiplication with no negative exponents in the answer, if so please help. (a^-2 c^1/5)^2/3

OpenStudy (anonymous):

start with \[a^{-2\times \frac{2}{3}}c^{\frac{1}{5}\times \frac{2}{3}}\]

OpenStudy (anonymous):

I have and got a^2/3 c 2/15 am I close?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

\[-2\times \frac{2}{3}=-\frac{4}{3}\]

OpenStudy (anonymous):

\[\large a^{-\frac{4}{3}}c^{\frac{2}{15}}\]

OpenStudy (anonymous):

Ok, so my answer is a^43/c^2/15

OpenStudy (anonymous):

Since the answer cannot have negative exponents I have to write it as a fraction correct?

OpenStudy (anonymous):

so I think the answer is a^4/3/c^2/15

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

writing with positive exponents does not mean to change the minus sign to a plus sign it means to put it in the denominator!

OpenStudy (anonymous):

\[\large \frac{c^{\frac{2}{15}}}{a^{\frac{4}{3}}}\]

OpenStudy (anonymous):

Ok, I think I understand now. I am being asked to write the answer in simplest form how do I do that? The writing is so small is that c^1/15 or c^2/15 over a ^4/2 ir ^4/3

OpenStudy (anonymous):

Never mind, I think I have it thank you!

OpenStudy (anonymous):

Here is your medal!